UT November 18, 2018

OSA: J. Williams

ISA: B. Rothberg (Tucson)

ISp: G. Bechetti

AO Support (FLAO): G. Taylor (Tucson)
AO Support (ARGOS): S. Walsh

Laser Dude: G. Rahmer

Software Support: A. Colson (Tucson)

Mtn. Manager. R. Hanson

UT 06:34 - started up LUCI software GUIs - will initialize and move mask out of LUCI-2 when at zenith for reconfig at 06:55

Handover negotiated to be at UT 07:10

UT 06:50 - initaizlaing LUCI-1

UT 07:06 - initialized LUCI-2 - putting 0.5" LS back into cabinet

UT 07:12 - configured for DX AROGS, SX SL LUCI

L1/L2 0001-0002 2.51 sec test darks

UT 07:16 - got co-pointing error on both sides - re-authorizing now

UT 07:21 - pointing check

UT 07:28 - collimation on bright star - SX is struggling

UT 07:33 - SX RIP'd - restarting

UT 07:40 - SX fixed, however, we are going to start with DX only just to make sure we can launch the lasers

UT 07:44 - seding monocular ARGOS collimation script to Pal2

UT 07:46 - test firing lasers -

UT 07:48 - collimation script OK - Step 1

UT 07:49 - sending real ARGOS script (this sends the AO Ref star for the AO system to acquire) - Step 2

UT 07:52 - AO ref star not visible - will have to spiral search with paddle

UT 07:53 - did pointing before switching to DX only. Wolfgang suggests that the miss means pointing was affected by the re-authorization - managed to find star

UT 07:57 - looks like we acquired a double star in the field, not the correct AO ref star

UT 07:59 - sent again, pointing looks good

UT 08:08 - new script made since AO Ref star shows as a double in CCD47. However, we used it last night with ESM with no problem

- made a new script (Palomar2_ARGOS_L2_NewAO.xml) - uses a fainter AO ref star above in DETY the old one

UT 08:11 - we are at 88 degree elevation - transiting - so can't fire

UT 08:20 - waiting 5 min for better elevation

UT 08:25 - resending new script

UT 08:26 - no star in CCD47

UT 08:32 - pointing check

UT 08:34 - paddle seems backwards - for moving star for pointing

UT 08:36 - resending NewAO Pal2 script

UT 08:37 - star appeared OK - have 30min on target

UT 08:40 - safe skip on secondary - inserted pause

UT 08:41 - sent script again - reacquiring

UT 08:42 - loops say closed taking first exposures - 0003-0004

-paused

UT 08:46 - lasers jumped?

UT 08:54 - laser propagating - not taking data - Gustavo troubleshooting vibration issues with Wolfgang

UT 09:01 - netowrk drop out - lost conneciton to OBS4 and polycom

UT 09:03 -back

UT 09:16 - d'plane d'plane! turning off laser

Important Note: After lasers are turned off we have to reacquire since telescope is in track mode (using lasers to guide on target)

UT 09:19 - Josh is running pointing checks - while we wait

UT 10:01 - IT 7463 - PCS shows a PA = 0 regardless of what the rotator is doing, this is affecting the paddle and may have other impacts - Josh is talking to Igor about it.

UT 10:04 - sending Palomar 2 DX only ARGOS collimation preset

UT 10:12 - sending ARGOS imaging preset

L2

0005 - loops closed but laser turned off first exposure

0006 - safe skip - loops closed on last seconds of this exposure

0007-008 - AO ref star is out of field but lasers are on

UT 10:23 - lost connection to mountain again

UT 10:24 - DX RIP'd - setting pause -

0013-0014 are bad - due to shell RIPd

UT 10:27 - shell RIP'd again while attempting to re-acquire AORef star

UT 10:30 - resending preset (new version with modified offset so AORef star should not be out of field)

UT 10:32 - clouds rolling back in - lots of closures coming, as soon as loops close the shell RIP'd

UT 10:36 - com dropout to mountain again!!!

UT 10:39 - right preset cancelled - Arbitrator failed - shell RIP'd

UT 10:45 - recovered

UT 10:47 - getting errors that cannot get rotator angle and elevation

-satellite software timestamp from TCS is stale by several minutes - TCS

UT 10:53 - tried reconnecting to TCS, failed, but finally reconnected - network issues?

UT 10:54 - same error as before cannot connect to - rotator and elevation angle

UT 10:58 - tried reconnect to IIF - DD says it connects, but still getting "could not get rotator and angle elevation"

AO can't connect to IIF either - bringing up telescope service via the LMC shows unknown state (the open_observer luci telsvc must be stale)

UT 11:10 -parking the telescope and closing the dome - dark thick clouds are overhead - we are unsure of TCS connection - erring on the side of caution

UT 11:14 - compensation mirror on LUCI-2 is stuck - stopped and started services

UT 11:16 - was able to reconnect to IIF
UT 11:40 - talking to Igor - we think connection is OK for LUCI-2, going to reauthorize the same config to see if this clears up the issue and ARGOS can communicate to the IIF

-whenever ARGOS tries to reconnect it fails because it gets an invalid number (the same large numbers that were in the facsum)

UT 11:43 - reauthoirzation cleared it up for ARGOS processes

UT 11:44 - going to send a test preset to see what happens

UT 11:47 - network drop out to mountain again!!!!
UT 11:55 - Josh and Igor testing the issues with pointing from earlier - clouds thick across the east (eliminating targets with RA>8hr)

UT 12:05 - officially ending laser operations

UT 12:18 - network drop out to mountain again!!!

UT 12:19 - closing up for the night....

-- BarryRothberg - 18 Nov 2018
Topic revision: r1 - 18 Nov 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback